The most recent financial statements for Moose Tours, Inc., appear below. Sales for 2016 are projected to grow by 20 percent. Interest expense will remain constant; the tax rate and the dividend payout rate will also remain constant. Costs, other expenses, current assets, fixed assets, and accounts payable increase spontaneously with sales.

MOOSE TOURS, INC.
2015 Income Statement
  Sales $ 745,000
  Costs 580,000
  Other expenses 16,000
  Earnings before interest and taxes $ 149,000
  Interest expense 12,000
  Taxable income $ 137,000
  Taxes (35%) 47,950
  Net income $ 89,050
  Dividends $ 21,920
  Addition to retained earnings 67,130

  

MOOSE TOURS, INC.
Balance Sheet as of December 31, 2015
Assets Liabilities and Owners’ Equity
  Current assets   Current liabilities
    Cash $ 20,440     Accounts payable $ 54,600
    Accounts receivable 32,760     Notes payable 13,800
    Inventory 69,720
      Total $ 68,400
      Total $ 122,920   Long-term debt $ 128,000
  Fixed assets   Owners’ equity
    Net plant and equipment $ 480,000     Common stock and paid-in surplus $ 114,000
    Retained earnings 292,520
      Total $ 406,520
  Total assets $ 602,920   Total liabilities and owners’ equity $ 602,920

   

What is the EFN if the firm was operating at only 80 percent of capacity in 2015? Assume that fixed assets are sold so that the company has a 100 percent asset utilization. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and round your answer to the nearest whole number, e.g., 32.)
Answer #1
Moose Tours,Inc.
Income Statement
2015 2016(20% increase in Sales) 20% increase in Sales
Sales=(A) $     7,45,000.00 745000*1.20 $                      8,94,000.00
Cost=(B) $     5,80,000.00 580000*1.20 $                      6,96,000.00
Other Expenses=(C ) $         16,000.00 16000*1.20 $                          19,200.00
Earnings before interest and tax=(D)=(A)-(B)-(C ) $     1,49,000.00 149000*1.20 $                      1,78,800.00
Interest Expense=(E ) $         12,000.00 same $                          12,000.00
Taxable Income=(F )=(D)-(E) $     1,37,000.00 $                      1,66,800.00
Taxes=(G) $         47,950.00 $                          58,380.00
Net Income=(H)=(F)-(G) $         89,050.00 $                      1,08,420.00
* Dividends=(I) $         21,920.00 $                          26,688.00
Addition to Retained Earnings=(H)-(I) $         67,130.00 $                          81,732.00
* Dividend=($21920/$89050)*$108420 $         26,688.00
Moose Tours,Inc.
Balance Sheet
2015 2016 2016
Assets
Cash $         20,440.00 20400*1.20 $                                24,528
Accounts Receivable $         32,760.00 32760*1.20 $                                39,312
Inventories $         69,720.00 69720*1.20 $                                83,664
Net Fixed Assets $     4,80,000.00 480000*1.20 $                            5,76,000
Total Assets(1) $     6,02,920.00 $                            7,23,504
Liabilties
Accounts Payable $         54,600.00 54600*1.20 $                                65,520
Notes Payables $         13,800.00 same $                                13,800
Long Term Debts $     1,28,000.00 same $                            1,28,000
Common Stock and paid in surplus $     1,14,000.00 same $                            1,14,000
** Retained Earnings $     2,92,520.00 $                            3,74,252
Total Liabilities & Equity(2) $     6,02,920.00 $                            6,95,572
AFN=(1)-(2) ($723504-$695572) $                                27,932
** Retained Earnings 31.12.2017
Retained Earnings (1.1.2017) $           2,92,520
Add: Net Income during 2017 $           1,08,420
Less: Dividend $               26,688
Retained Earnings $           3,74,252